Two 6µF capacitance are connected in series with one having a capacitance of 12µF.Find the total equivalent circuit capacitance. What capacitance must be added in series to obtain a capacitance of 1.2µF.
1
Expert's answer
2020-05-07T08:40:48-0400

Explanations

  • The equivalent capacitance of a set of capacitors connected in series is given by

"\qquad\qquad
\small \frac{1}{C_{eq}} = \small \frac{1}{C_1}+ \frac{1}{C_2}+ \frac{1}{C_3} + \cdots\\"

  • Equivalent of a set connected in parallel is given by

"\qquad\qquad
\small C_{eq} = \small C_1 + C_2 + C_3 + \cdots"


Calculations


1) Since all the capacitors are connected in series,

"\qquad 
\begin{aligned}
\small \frac{1}{C_{eq}} &= \small \frac{1}{6\mu F}+ \frac{1}{6\mu F}+ \frac{1}{12\mu F} \\
\small C_{eq} &= \small \bold{2.4\mu F}
\end{aligned}"


2)

If the needed capacitance is C then,

"\qquad
\begin{aligned}
\small \frac{1}{1.2} &= \small \frac{1}{2.4\mu F}+ \frac{1}{C}\\
\small C &= \small \bold{2.4\mu F}
\end{aligned}"